# Rambler: Supporting Writing With Speech via LLM-Assisted Macro Revision

### 19 Jan 2024 v1

@@ 3.3. Example User Workflow
To better illustrate the functionality of Rambler, we consider an example of a simple user workflow: Alex, an artist, is drafting a post for his personal blog and wants to describe how he finds creative inspiration. He has many ideas on the topic, but isn’t sure the best way to write it: hence, he reasons that he should simply talk it out. Opening up Rambler, Alex presses the + Ramble button to begin recording. After finishing his thoughts and stopping the recording, Rambler automatically cleans this into the first Ramble. Alex senses that he wandered somewhat from his initial topic; since the resulting text is fairly long, he slides the semantic zoom slider to see a summarized version of his text. Then he takes a look at the highlighted keywords in all the zoom levels to get a better sense of what he talked about. After glancing back at the transcript, Alex decides to utilize semantic splitting to revise the various ideas in his Ramble separately.
Rambler segments his cleaned spoken text into multiple rambles along three main ideas, 1) ”finding personal inspiration”, 2) ”interaction between self expression and audience interpretation”, and 3) ”difference between art and design.” With a clearer understanding of the ideas he’s just verbalized, Alex begins to alternate between composing new related Rambles, and revising with the provided functionalities (eg. respeaking, reordering, merging, and/or splitting Rambles) – until he feels like each Ramble roughly resembles a somewhat concrete idea. Once he feels that he has added all his thoughts and developed them into a cohesive flow of ideas, he uses the built-in keyboard to fix word choices here and there to his liking, and then exports the composition to share with the world.
@@ 3.4. Prototype Implementation
Rambler is implemented as a React.js web application with a node.js backend and is hosted in the cloud. LLM-based text manipulation functionality is enabled using OpenAI’s GPT-4 API, which provides high quality output for our desired LLM functionality. We considered other models such as GPT-3.5 Turbo for response time reasons; their output on LLM interactions (most notably, semantic split) was less precise. Rambler’s voice transcription uses AssemblyAI’s real time Speech-to-Text API. We made this engineering choice because AssemblyAI’s API allowed us to transcribe text indefinitely in a single instance. Automatic keyword selection is implemented via the rapid automatic keyword extraction (RAKE) method (Rose et al., 2010). Compared to an LLM based approach for keyword selection akin the method used in Graphologue (Jiang et al., 2023), the RAKE method was orders of magnitude faster ($\sim$4s vs nearly instant). Our overall system architecture is illustrated with Figure 6. Our project source code is available at [LINK ANONYMIZED] and a demo can be found at [LINK ANONYMIZED].
GPT-4 was used for the four key LLM-powered gist-related features: text summarization for semantic zoom (including transcription cleaning), merging multiple Rambles, splitting a single Ramble into multiple, and open-ended editing using customized LLM prompts. We created pre-defined prompts for zoom, merging, and splitting; the open-ended editing prompts are entirely user defined. One-shot prompting was employed for the transcription cleaning prompt and zero-shot was used for the rest. Keywords are appended to prompts when marked for inclusion through selection. All prompts are attached in Appendix A.2 for reference.
@@ 4. Evaluation
To evaluate the effectiveness of Rambler for supporting writing with speech, we designed a within-subjects study to compare Rambler with a state-of-the-art solution for dictation and word processing. The baseline is chosen to be: raw dictation into a text editor with ChatGPT provided on the side. Participants completed two long-form writing tasks through dictation, one using Rambler’s full suite of available functionality and the other using the baseline. Through this comparison, we seek answers to the following questions:
@@
    Can writing with speech be better supported by Rambler than the baseline? If so, how?
@@
    How effectively does our approach of embedding LLMs in GUI support the review and revision of spoken text?
@@ 4.1. Baseline Interface
As mentioned above, the baseline uses a plain text editor interface plus dictation. Raw transcripts go into the text box as a user speaks after the microphone is activated. The transcript can be edited using iOS’s built-in keyboard and cursor interface. To avoid noises caused by different speech recognition engines, we implemented this interface using the same speech recognition engine (Assembly.ai) used in our prototype as a web interface. To make a fair comparison between the baseline and Rambler, ChatGPT is provided with a separate tab in the same web browser. However, we limited the use of ChatGPT to revising content produced by the participants by not allowing them to ask ChatGPT to produce major compositions for them.
@@ 4.2. Participants
Twelve participants ($N=12$) were recruited from our university campus community through various mailing lists and student organizations. All participants were native-fluent English speakers and there was varied experience with dictation and LLM use. P3 and P10 were aged between 25 and 34; all other participants were aged between 18 and 24. No participant had motor impairments that impacted keyboard or dictation usage. Further participant details are presented in Table 1.
@@ 4.3. Experiment Task
The task for the experiment is to compose and revise a blog post using the given interface in each condition. The reason for this task setting is to encourage participants to compose long-form personal writing and revise it extensively for an audience, in order to combat the potential lack of motivation to improve the text in a lab study. Participants were asked to choose a topic they were personally interested in, from a provided list of open-ended writing tasks (e.g., “Think of two dream jobs of yours.”, “Think of two academic concepts you find interesting.”, and so on). Then, for each of the two conditions, they were asked to create a “blog post that would be posted on your social media” about one chosen answer to the topic. Notice that we had participants pick two variants for the same topic (such as two dream jobs instead of a dream job and an academic concept) to minimize the influence that completely different topics would have on participants’ compositions.
@@ 4.4. Procedure
A lab study session for each participant lasted about 90 minutes, which could be divided into three main parts – the pre-study survey, the main study, and the post-study interview. The study session was conducted in a quiet room in person with an experimenter. Participants accessed Rambler via Safari on an 11-inch iPad Pro running iOS 16. There was no hardware keyboard. Nonetheless, the participants were allowed to use the on-screen keyboard when needed.
The pre-study survey.
@@
The main study.
Participants were not told we made the interfaces. In a counterbalanced order across participants, each participant used both two conditions. For each condition, the participants engaged in 1) a training trial, 2) the task, and 3) a post-task survey. In the training trial, the interviewer spent about 5 to 15 minutes explaining all the functionality of the given condition and provided time for the participant to try and familiarize themselves with each feature. Next, to perform the task, the participants were given up to 20 minutes to improve their composition until the time is up, or until they felt satisfied, whichever came first. If the participants hit a word count of 800 words or more, or had not done any editing by the 10-minute mark, the experimenter would encourage the participant to begin editing (so that they have a decent amount of time to revise their composition). In the post-task survey, the participants were asked to assess the usefulness of each feature in the given condition using a five-point Likert scale. They were also asked to estimate a percentage on how much micro-revision (specifically keyboard-editing) they did in comparison to macro-revisions. They were then asked to assess the quality of their text composition on a five-point Likert scale.
Table 1. Demography and dictation/LLM usage of study participants.
ID Degree Level Domain Dictation Usage LLM Usage
@@
Computer Science
A few times a year
A few times a week
@@
Computer Science
A few times a month
A few times a week
@@
A few times a week
A few times a month
@@
A few times a month
A few times a month
@@
Applied Physics
A few times a year
A few times a month
@@
The post-study interview.
Following the two conditions, the participants were interviewed for about 25 minutes regarding their overall experience. Specifically, they were asked about their workflows for each condition and their experience with the features in both conditions. Finally, the participants were asked for their overall preference between the baseline and the Rambler condition. Each participant was compensated \$30 USD for the study.

### 08 Mar 2024 v2

@@ 3.3. Example User Workflow
To better illustrate the functionality of Rambler, consider the following example of a simple user workflow: Alex, an artist, is drafting a post for his personal blog and wants to describe how he finds creative inspiration. He has many ideas on the topic, but isn’t sure the best way to write it: hence, he reasons that he should simply talk it out. Opening up Rambler, Alex presses the + Ramble button to begin recording. After finishing his thoughts and stopping the recording, Rambler automatically cleans his dictated text into the first Ramble. Alex senses that he wandered somewhat from his initial topic; since the resulting text is fairly long, he slides the Semantic Zoom slider to see a summarized version of his text. Then he takes a look at the highlighted keywords in all the zoom levels to get a better sense of what he talked about. After glancing back at the transcript, Alex decides to use Semantic Split to revise the various ideas in his Ramble separately.
Rambler segments his cleaned spoken text into multiple rambles along three main ideas, 1) “finding personal inspiration,” 2) “interaction between self expression and audience interpretation,” and 3) “difference between art and design.” With a clearer understanding of the ideas he’s just verbalized, Alex begins to alternate between composing new Rambles, and revising with the provided functionalities (e.g., respeaking, reordering, merging, and/or splitting Rambles) – until he feels like each Ramble roughly resembles a somewhat concrete idea. Once he feels that he has added all his thoughts and developed them into a cohesive flow of ideas, he uses the built-in onscreen keyboard to fix word choices here and there to his liking, and then exports his composition to share with the world.
Figure 6. Rambler System Architecture Diagram consisting of a web-based frontend, a cloud server mediating between user requests and the OpenAI API for LLM functionality. The AssemblyAI API is used for real-time speech transcription.
@@ 3.4. Prototype Implementation
Rambler is implemented as a React.js web application with a node.js backend and is hosted in the cloud. LLM-based text manipulation functionality is enabled using OpenAI’s GPT-4 API, which provides high quality output for our desired LLM functionality. We considered other models such as GPT-3.5 Turbo for response time reasons; their output on LLM interactions (most notably, Semantic Split) was less precise. Rambler’s voice transcription uses AssemblyAI’s real time Speech-to-Text API. We made this engineering choice because AssemblyAI’s API allowed us to transcribe text indefinitely in a single instance. Automatic keyword selection is implemented via the rapid automatic keyword extraction (RAKE) method (Rose et al., 2010). Compared to an LLM based approach for keyword selection akin the method used in Graphologue (Jiang et al., 2023), the RAKE method was orders of magnitude faster ($\sim$4s vs nearly instant). Our overall system architecture is illustrated with Figure 6. Our project source code is available at https://github.com/BerkeleyHCI/rambler.
We used GPT-4 to power the four key LLM-based gist-related features: text summarization for Semantic Zoom (including transcription cleaning), merging multiple Rambles, splitting a single Ramble into multiple, and open-ended editing using customized LLM prompts. We created pre-defined prompts for zoom, merging, and splitting; the open-ended editing prompts are entirely user defined. One-shot prompting was employed for the transcription cleaning prompt and zero-shot was used for the rest. Keywords are appended to prompts when marked for inclusion through selection. All prompts are attached in Appendix A.2 for reference.
@@ 4. Evaluation
To evaluate the effectiveness of Rambler for supporting writing with speech, we designed a within-subjects study to compare Rambler with a state-of-the-art solution for dictation and word processing. We chose, as our baseline, raw dictation into a text editor with ChatGPT provided on the side. Participants completed two long-form writing tasks through dictation, one using Rambler’s full suite of available functionality, and the other using the baseline. Through this comparison, we seek answers to the following questions:
@@
    Is writing with speech better supported by Rambler than the baseline? If so, how?
@@
    How effectively does our approach of embedding LLMs in a GUI support the review and revision of spoken text?
@@ 4.1. Baseline Interface
As mentioned above, the baseline uses a plain text editor interface plus dictation. Raw transcripts populate the text content area as a user speaks, after the microphone is activated. The transcript can be edited using iOS’s built-in keyboard and cursor interface. To avoid potential confounding effects resulting from differing speech recognition engines, we implemented our baseline interface using the same speech recognition engine used in our prototype (Assembly.ai), as a web interface. For a fair comparison between the baseline and Rambler, in our baseline condition we provided ChatGPT through a separate tab in the same web browser—thus ensuring that our results are primarily driven by the affordances of Rambler rather than the availability of an LLM. However, we did limit the use of ChatGPT to revising content produced by participants, by discouraging participants from asking ChatGPT to generate lengthy compositions.
@@ 4.2. Participants
Twelve participants ($N=12$), seven who self-reported as female and five as male, were recruited from our university campus community through various mailing lists and student organizations. All participants were native-fluent English speakers, and there was varied experience with dictation and LLM use. P3 and P10 were aged between 25 and 34; all other participants were aged between 18 and 24. No participant had motor impairments that impacted keyboard or dictation usage. Participant details are presented in Table 1.
@@ 4.3. Experiment Task
Participants were asked to compose and revise a blog post using the given interface in each condition. We chose this task to encourage participants to compose long-form personal writing and revise it extensively for an audience, in order to address a potential lack of motivation, in a lab study setting, to improve text the participants may not have any reason to feel invested in. Thus participants were asked to choose a topic they were personally interested in, from a provided list of open-ended writing tasks (e.g., “Think of two dream jobs of yours.” and “Think of two academic concepts you find interesting.”). For each of the two conditions, participants were asked to create a “blog post that would be posted on your social media” about one chosen answer to the topic. Note that we had participants use the same topic across conditions (e.g., two dream jobs, not one dream job and one academic concept), to mitigate the potential impact of topic selection on participants’ compositions.
@@ 4.4. Procedure
A lab study session for each participant lasted about 90 minutes, divided into three parts: the pre-study survey, the main study itself, and the post-study interview. The study session was conducted in a quiet room, in person, with an experimenter. Participants accessed Rambler via Safari on a researcher-provided 11-inch iPad Pro running iOS 16. There was no hardware keyboard, however, participants were allowed to use the on-screen keyboard as needed.
Pre-study survey.
@@
Main study.
Participants were not told we made the interfaces. In a counterbalanced order across participants, each participant used both interfaces in two conditions. For each condition, the participants engaged in 1) a tutorial for training, 2) the task, and 3) a post-task survey. In the training trial, the interviewer spent about 5 to 15 minutes explaining all the functionality of the given condition and provided time for the participant to try and familiarize themselves with each feature. Next, to perform the task, the participants were given up to 20 minutes to improve their composition until the time is up, or until they felt satisfied, whichever came first. If the participants hit a word count of 800 words or more, or had not done any editing by the 10-minute mark, the experimenter would encourage the participant to begin editing (to ensure enough time for revisions). In the post-task survey, the participants were asked to assess the usefulness of each feature in the given condition using a five-point Likert scale. They were also asked to estimate what percentage of their time was spent in micro-revision (specifically keyboard-editing), compared with macro-revision. Finally, participants were asked to assess the quality of their text composition on a five-point Likert scale.
Table 1. Demography and dictation + LLM usage of study participants.
ID Degree Domain Dictation Usage LLM Usage
@@
Comp. Sci.
Yearly
Weekly
@@
Comp. Sci.
Monthly
Weekly
@@
Weekly
Monthly
@@
Monthly
Monthly
@@
Physics
Yearly
Monthly
@@
Post-study interview.
Following the two conditions, participants were interviewed regarding their overall experience, for around 25 minutes. Specifically, they were asked about their workflows for each condition and their experiences with the features in both conditions. Finally, participants were asked for their overall preference between the baseline and the Rambler condition. Each participant was compensated \$30 USD for their participation.
@@ 4.5. Data Collection and Analysis
Additionally, we evaluated participant output text for fluency and grammaticality using the GRUEN automated benchmark (Zhu and Bhat, 2020), and compared our participants’ final text with similar “blog post” text sampled from a popular corpus of weblog posts (Schler et al., 2006) to match participant demographics on age and gender.